OpenStudy (anonymous):

What is the domain of this function? F(x) = 7/(6+ln(7x)) In interval notation?

OpenStudy (campbell_st):

well you can't take the log of a negative number or zero ... so x > 0 just work out the interval notation

OpenStudy (campbell_st):

but you also need to consider that \[6 + \ln(7x) \neq 0\] so this may help to find the domain so \[\ln (7x) = -6\] then \[7x = e^{-6}\] \[x = \frac{e^{-6}}{7} \] so that looks like the lower limit of the interval...

OpenStudy (anonymous):

Yeah, solved the denominator for zero and knew that limit, but forgot about it not being less than zero. Thanks!